# Low-Rank Graph Contrastive Learning for Node Classification

### v1

@@ C.2 Node Classification Results for GCL Methods with different types of classifiers
LR-GCL 0.858$\pm$0.006 0.589$\pm$0.011 0.713$\pm$0.007 0.695$\pm$0.011 0.492$\pm$0.011 0.587$\pm$0.013 0.477$\pm$0.012 0.306$\pm$0.012 0.419$\pm$0.012 0.363$\pm$0.011
C.3 Training Time Comparisons and Study on Different Ranks
In this section, we first compare the training time of LR-GCL against other baseline methods on all benchmark datasets. For our method, we include the training time of robust graph contrastive learning, the time of the SVD computation of the kernel, and the training time of the transductive classifier. For graph contrastive learning methods, we include both the training time of the GCL encoder and the downstream classifier. We evaluate the training time on one 80 GB A100 GPU. The results are shown in Table 7. It is observed that the LR-GCL takes similar training time as baseline GCL methods such as SFA and MERIT.
We also perform ablation study on the value of rank $r$ in the optimization problem (3) for our low-rank transductive classifier. It is observed from Table 2 that the performance of our low-rank classifier is consistently close to the best performance among all the choices of the rank when $r$ is between $0.1\min\left\{N,d\right\}$ and $0.2\min\left\{N,d\right\}$.
GCN 11.5 13.7 38.6 43.2 22.3 30.2 19.0 215.1
@@ C.3 Training Time Comparisons and Study on Different Ranks
GCE 32.6 36.9 67.3 80.8 37.6 50.1 32.2 346.1
UnionNET 67.5 69.7 100.5 124.2 53.2 69.2 45.3 479.3
NRGNN 72.4 80.5 142.7 189.4 74.3 97.2 62.4 650.2
RTGNN 143.3 169.5 299.5 353.5 153.7 201.5 124.2 1322.2
SUGRL 100.3 122.1 207.4 227.1 107.7 142.8 87.7 946.8
MERIT 167.2 179.2 336.7 375.3 172.3 226.5 140.6 1495.1
ARIEL 156.9 164.3 284.3 332.6 145.1 190.4 118.3 1261.4
SFA 237.5 269.4 457.1 492.3 233.5 304.5 187.2 2013.1
Sel-Cl 177.3 189.9 313.5 352.5 161.7 211.1 130.9 1401.1
Jo-SRC 148.2 157.1 281.0 306.1 144.5 188.0 118.5 1256.0
GRAND+ 57.4 68.4 101.7 124.2 54.8 73.8 44.5 479.2
LR-GCL 159.9 174.5 350.7 380.9 180.3 235.7 145.5 1552.7
C.4 Eigen-Projection and Concentration Entropy
Figure 3 illustrates the eigen-projection and energy concentration ratio for more datasets.
C.5 Visualization of Confidence Score
We visualize the confident nodes selected by BPL in the embedding space of the learned node representations in Figure 4. The node representations are visualized by the t-SNE figure. Each mark in t-SNE represents the representation of a node, and the color of the mark denotes the confidence of that node. The results are shown for different levels of attribute noise. It is observed from Figure 4 that confident nodes, which are redder in Figure 4, are well separated in the embedding space. With a higher level of attribute noise, the bluer nodes from different clusters blended around the cluster boundaries. In contrast, the redder nodes are still well separated and far away from cluster boundaries, which leads to more robustness and better performance in downstream tasks.
(a) noise level = 0
(b) noise level = 20
(c) noise level = 40
Figure 4: Visualization of confident nodes with different levels of attribute noise for semi-supervised node classification on Citeseer.
C.6 Study in the Kernel Complexity
With feature matrix $\mathbf{H}$ produced by different encoders, we compute a kernel matrix $\mathbf{K}=\widehat{}\mathbf{A}\mathbf{H}\mathbf{H}^{\top}\widehat{}\mathbf{A}^{\top}$ before and after optimization of problem (3). Score of $\mathbf{K}$ is
complexity of $\textup{complexity of }\mathbf{K}=\min_{h\in[0,n]}\frac{h}{n}+\sqrt{\frac{\sum\limits_{i=h+1}^{n}\widehat{\lambda}_{i}}{n}}$ (11)
Also, compute $\mathbf{K}$ for LR-GCL and other GCL baselines.
Datasets MERIT SFA RGCL LR-GCL
Cora Complex. 0.5485 0.5749 0.5383 0.1675
$h$ 1455 1532 1431 433 (rank)
Citeseer Complex. 0.4233 0.4546 0.4090 0.1178
$h$ 1347 1479 1330 366
PubMed Complex. 0.9770 0.1095 0.0932 0.0674
$h$ 1690 1874 1667 1183
Wiki-CS Complex. 0.1615 0.1762 0.1498 0.1062
$h$ 1784 1986 1639 1170
Amazon-Computers Complex. 0.1132 0.1217 0.1110 0.0655
$h$ 1467 1579 1433 874
Amazon-Photos Complex. 0.1875 0.2047 0.1823 0.1007
$h$ 1397 1532 1353 750
Coauthor-CS Complex. 0.0764 0.0859 0.0749 0.0622
$h$ 1657 1874 1605 1120
ogbn-arxiv Complex. 0.0124 0.0139 0.0117 0.0103
$h$ 1769 1920 1653 1354
Table 8: Comparisons in complexity of kernels.

### v2

@@ 5.10 Node Classification Results for GCL Methods with Different Types of Classifiers
GCL-LRR 0.858$\pm$0.006 0.589$\pm$0.011 0.713$\pm$0.007 0.695$\pm$0.011 0.492$\pm$0.011 0.587$\pm$0.013 0.477$\pm$0.012 0.306$\pm$0.012 0.419$\pm$0.012 0.363$\pm$0.011
GCL-LR-Attention 0.861$\pm$0.006 0.602$\pm$0.011 0.724$\pm$0.007 0.708$\pm$0.011 0.510$\pm$0.011 0.605$\pm$0.013 0.492$\pm$0.012 0.329$\pm$0.012 0.436$\pm$0.012 0.382$\pm$0.011
5.11 Statistical Significance Analysis
In this section, we compute the p-values from paired t-tests comparing the performance of GCL-LRR and GCL-LR-Attention against the strongest baseline methods to assess the statistical significance of the observed improvements. As shown in Table 11, the p-values for both GCL-LRR and GCL-LR-Attention remain consistently below the threshold of $0.05$ across all datasets and noise conditions, thereby confirming that the performance gains over the top baselines are statistically significant.
Table 11: P-values of the t-tests for GCL-LRR and GCL-LR-Attention against the top baseline methods under each noise setting on all the benchmark datasets.
Datasets Methods Noise Type
40 60 80
Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute
Cora GCL-LRR 0.038 0.024 0.021 0.035 0.021 0.041 0.025 0.028 0.031
GCL-LR-Attention 0.027 0.022 0.018 0.018 0.038 0.035 0.031 0.023 0.030
Citeseer GCL-LRR 0.043 0.035 0.022 0.021 0.030 0.041 0.027 0.024 0.022
GCL-LR-Attention 0.037 0.038 0.019 0.027 0.025 0.040 0.035 0.037 0.030
PubMed GCL-LRR 0.028 0.043 0.030 0.026 0.027 0.043 0.036 0.040 0.042
GCL-LR-Attention 0.025 0.030 0.026 0.023 0.024 0.041 0.033 0.035 0.037
Coauthor-CS GCL-LRR 0.041 0.032 0.036 0.043 0.044 0.040 0.027 0.037 0.042
GCL-LR-Attention 0.036 0.030 0.034 0.041 0.042 0.039 0.025 0.033 0.036
ogbn-arxiv GCL-LRR 0.040 0.032 0.036 0.043 0.044 0.040 0.027 0.037 0.042
GCL-LR-Attention 0.036 0.030 0.034 0.041 0.042 0.039 0.025 0.033 0.036
Wiki-CS GCL-LRR 0.044 0.035 0.033 0.028 0.034 0.041 0.036 0.040 0.042
GCL-LR-Attention 0.040 0.036 0.031 0.026 0.029 0.039 0.034 0.038 0.040
Amazon-Computers GCL-LRR 0.033 0.043 0.034 0.031 0.034 0.041 0.036 0.040 0.042
GCL-LR-Attention 0.031 0.038 0.030 0.028 0.030 0.038 0.034 0.037 0.040
Amazon-Photos GCL-LRR 0.040 0.041 0.038 0.043 0.044 0.040 0.027 0.037 0.042
GCL-LR-Attention 0.037 0.039 0.036 0.041 0.042 0.039 0.025 0.033 0.036
Texas GCL-LRR 0.038 0.024 0.021 0.035 0.021 0.041 0.025 0.028 0.031
GCL-LR-Attention 0.027 0.022 0.018 0.018 0.038 0.035 0.031 0.023 0.030
Chameleon GCL-LRR 0.044 0.035 0.033 0.028 0.034 0.041 0.036 0.040 0.042
GCL-LR-Attention 0.040 0.036 0.031 0.026 0.029 0.039 0.034 0.038 0.040
Table 12: Sensitivity analysis on the weighting parameter $\tau$ for the TNN. The study is performed using GCL-LR-Attention on the Coauthor-CS dataset for semi-supervised node classification under $60\%$ asymmetric label noise.
$\tau$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Accuracy 0.588 0.590 0.593 0.591 0.594 0.590 0.591 0.590 0.589
5.12 Sensitivity Analysis on the Hyperparameters
We perform a sensitivity analysis on the weighting parameter $\tau$ associated with the TNN ${\left\|\mathbf{K}\right\|}_{r_{0}}$ in Equation 1. This analysis is conducted using GCL-LR-Attention on the Coauthor-CS dataset under the setting of semi-supervised node classification with $60\%$ asymmetric label noise. We vary $\tau$ over the set $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$, and the corresponding classification accuracies are reported in Table 12. The highest accuracy is achieved at $\tau=0.5$, though GCL-LR-Attention maintains stable and competitive performance across the entire range. Even in the least favorable case, with $\tau=0.1$, the accuracy declines by only $0.6\%$, highlighting the robustness of the model to variations in $\tau$.
6 Conclusions
This paper first introduces a novel graph contrastive learning encoder, Graph Contrastive Learning with Low-Rank Regularization (GCL-LRR), designed to enhance the robustness in node representation learning. GCL-LRR generates low-rank features, drawing motivation from the low-frequency characteristics commonly observed in real-world graph datasets and the tight generalization bound for transductive learning. The GCL-LRR encoder is trained under a prototypical GCL framework, incorporating the TNN as a regularization term. We assess the effectiveness of GCL-LRR through extensive comparisons with existing methods on semi-supervised and transductive node classification tasks, where the input graphs are perturbed by either label noise or attribute noise. To further strengthen its performance, we extend GCL-LRR with a novel low-rank attention mechanism, leading to the GCL-LR-Attention model. GCL-LR-Attention further reduces the kernel complexity of GCL-LRR, thereby improving the generalization of GCL-LRR. Empirical results across a wide range of benchmarks confirm that both GCL-LRR and GCL-LR-Attention achieve superior robustness and outperform state-of-the-art approaches in learning effective node representations for noisy node classification tasks.
7 Theoretical Results
We present the proof of Theorem 4.1 in this section.
Proof of Theorem 4.1.
Define $\mathbf{N}\coloneqq\mathbf{Y}-\tilde{}\mathbf{Y}\in\mathbb{R}^{N}$ as the label noise. It can be verified that at the $t$-th iteration of gradient descent for $t\geq 1$, we have
$\mathbf{W}^{(t)}$ $=\mathbf{W}^{(t-1)}-\eta\left[\mathbf{H}\right]_{{\cal L}}^{\top}\left[\mathbf{H}\mathbf{W}^{(t-1)}-\mathbf{Y}\right]_{{\cal L}}$
$=\mathbf{W}^{(t-1)}-\eta\left[\mathbf{H}\right]_{{\cal L}}^{\top}\left[\mathbf{H}\mathbf{W}^{(t-1)}-\tilde{}\mathbf{Y}\right]_{{\cal L}}+\eta\left[\mathbf{H}\right]_{{\cal L}}^{\top}\left[\mathbf{N}\right]_{{\cal L}}.$ (5)
It follows by (7) that
$\left[\mathbf{H}\right]_{{\cal L}}\mathbf{W}^{(t)}=\left[\mathbf{H}\right]_{{\cal L}}\mathbf{W}^{(t-1)}-\eta{\mathbf{K}}_{{\cal L},{\cal L}}\left[\mathbf{H}\mathbf{W}^{(t-1)}-\tilde{}\mathbf{Y}\right]_{{\cal L}}+\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\left[\mathbf{N}\right]_{{\cal L}},$ (6)
where $\mathbf{K}_{{\cal L},{\cal L}}\coloneqq\left[\mathbf{H}\right]_{{\cal L}}\left[\mathbf{H}\right]_{{\cal L}}^{\top}\in\mathbb{R}^{m\times m}$. With $\mathbf{F}(\mathbf{W},t)=\mathbf{H}\mathbf{W}^{(t)}$, it follows by (6) that
$\left[\mathbf{F}(\mathbf{W},t)-\tilde{}\mathbf{Y}\right]_{{\cal L}}=\left(\mathbf{I}_{m}-\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\right)\left[\mathbf{F}(\mathbf{W},t-1)-\tilde{}\mathbf{Y}\right]_{{\cal L}}+\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\left[\mathbf{N}\right]_{{\cal L}}.$
It follows from the above equality and the recursion that
$\left[\mathbf{F}(\mathbf{W},t)-\tilde{}\mathbf{Y}\right]_{{\cal L}}=-\left(\mathbf{I}_{m}-\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\right)^{t}\left[\tilde{}\mathbf{Y}\right]_{{\cal L}}+\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\sum_{t^{\prime}=0}^{t-1}\left(\mathbf{I}_{m}-\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\right)^{t^{\prime}}\left[\mathbf{N}\right]_{{\cal L}}$ (7)
We apply [73, Corollary 3.7] to obtain the following bound for the test loss $\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\tilde{}\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$:
$\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\tilde{}\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$ $\leq\frac{c_{0}}{m}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\tilde{}\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}^{2}+c_{0}\min_{0\leq Q\leq N}r(u,m,Q)+\frac{c_{0}x}{u},$ (8)
with
$r(u,m,Q)\coloneqq Q\left(\frac{1}{u}+\frac{1}{m}\right)+\left(\sqrt{\frac{\sum\limits_{q=Q+1}^{N}\widehat{\lambda}_{q}}{u}}+\sqrt{\frac{\sum\limits_{q=Q+1}^{N}\widehat{\lambda}_{q}}{m}}\right),$
where $c_{0}$ is a positive constant depending on $\mathbf{U}$, $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{r}$, and $\tau_{0}$ with $\tau_{0}^{2}=\max_{i\in[N]}\mathbf{K}_{ii}$.
It follows from (7) and (8) that for every $r_{0}\in[N]$, we have
$\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\tilde{}\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}\leq\frac{c_{0}}{m}{\left\|\left(\mathbf{I}_{m}-\eta\left[\mathbf{K}\right]_{{\cal L},{\cal L}}\right)^{t}\left[\tilde{}\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}^{2}$
$\hskip 93.95122pt+c_{0}r_{0}\left(\frac{1}{u}+\frac{1}{m}\right)+c_{0}\left(\sqrt{\frac{\sum\limits_{q=r_{0}+1}^{N}\widehat{\lambda}_{q}}{u}}+\sqrt{\frac{\sum\limits_{q=r_{0}+1}^{N}\widehat{\lambda}_{q}}{m}}\right)+\frac{c_{0}x}{u}$
@@ 7 Theoretical Results
$\phantom{=}+c_{0}r_{0}\left(\frac{1}{u}+\frac{1}{m}\right)+c_{0}\sqrt{{\left\|\mathbf{K}\right\|}_{r_{0}}}\left(\sqrt{\frac{1}{u}}+\sqrt{\frac{1}{m}}\right)+\frac{c_{0}x}{u},$ (9)
[…]